OpenStudy (anonymous):

can i get some help? i'm a little lost. Suppose that you are climbing a hill whose shape is given by z=833−0.05x^2−0.02y^2, and that you are at the point (90, 80, 300). In which direction (unit vector) should you proceed initially in order to reach the top of the hill fastest? ⟨ , ⟩ If you climb in that direction, at what angle above the horizontal will you be climbing initially (radian measure)?

OpenStudy (anonymous):

i thought it would just be gradient = <-.1x, -.04y> making part a be <9,3.2>, but that wasn't right.

OpenStudy (aum):

gradient = <-.1x, -.04y> at x = 90, y = 80: gradient = <-9, -3.2> "In which direction (UNIT VECTOR) should you proceed ...." | gradient | = sqrt( (-9)^2 + (-3.2)^2 ) = ? unit vector = (-9i - 3.2j) / | gradient | = ?

OpenStudy (aum):

I am getting -0.9422i - 0.3350j

OpenStudy (anonymous):

oh. i never know when i have to divide by the magnitude and when i don't.....

OpenStudy (aum):

If they ask for UNIT VECTOR you will have to divide by the magnitude. Only then the magnitude of the vector will be 1.

OpenStudy (anonymous):

so how do i find the degree? part b

OpenStudy (anonymous):

i know Duf= gradf *u *costheta, but not sure what to use in that equation.

OpenStudy (aum):

The initial position vector is <90, 80, 300>. Find the magnitude. The gradient vector is <-9, -3.2, 0>. Find the magnitude. cos(theta) = | gradient | / | position | set calculator to radian mode and take the inverse since they want the angle in radians.

OpenStudy (anonymous):

put this into webwork....arccos(sqrt((-9)^2+(-3.2)^2)/sqrt(90^2+80^2+300^2)) and it says incorrect....

OpenStudy (aum):

Is the mode set to radians?

OpenStudy (anonymous):

in webwork, you can't change degrees or radians. it always does it in radians.

OpenStudy (aum):

I am getting 1.5412 radians. Can you try that?

OpenStudy (anonymous):

1.54124358964091

OpenStudy (aum):

Is it saying that is wrong? Did it accept answer to part a) ?

OpenStudy (anonymous):

getting a screenshot. two secs.

OpenStudy (anonymous):

ok, i'll check with the teacher. thanks

OpenStudy (aum):

How many tries are you allowed?

OpenStudy (anonymous):

unlimited

OpenStudy (aum):

Wondering if sticking a pi - in front would work.

OpenStudy (aum):

pi minus arccos(....)

OpenStudy (anonymous):

nope

OpenStudy (aum):

Alright I'll leave it to your teacher.

OpenStudy (anonymous):

thanks!

OpenStudy (aum):

you are welcome. If you do find out the answer you can post it here and I'll check it the next time.

OpenStudy (aum):

The gradient vector is <-9, -3.2> @saiken2009 Could you try arctan( (-9)^2 + (-3.2)^2 ) and tell me if it accepts the answer?

OpenStudy (anonymous):

no, it doesn't.